# Explaining time series models using frequency masking

### v1

@@ Abstract
Time series data is fundamentally important for describing many critical domains such as healthcare, finance, and climate, where explainable models are necessary for safe automated decision-making. To develop eXplainable AI (XAI) in these domains therefore implies explaining salient information in the time series. Current methods for obtaining saliency maps assumes localized information in the raw input space. In this paper, we argue that the salient information of a number of time series is more likely to be localized in the frequency domain. We propose FreqRISE, which uses masking based methods to produce explanations in the frequency and time-frequency domain, which shows the best performance across a number of tasks.
Machine Learning, ICML
@@ 1 Introduction
With the increasing development of systems for automated decision-making based on time series in critical domains such as healthcare (Phan & Mikkelsen, 2022; Brüsch et al., 2023), finance (Giudici & Raffinetti, 2023) and climate forecasting (González‐Abad et al., 2023), the demand for safe and trustworthy machine learning models is ever rising. High accuracy and explainability are both necessary components in achieving safety and trustworthiness. Deep learning models have become a popular choice to obtain high accuracy for time series tasks (Mohammadi Foumani et al., 2024), but due to their complex reasoning process they are difficult to interpret. Explainable artificial intelligence (XAI) aims to open up the black box (Samek et al., 2019).
@@
Most advancements in the field of XAI have been made in explaining image inputs (Kim et al., 2018; Petsiuk et al., 2019; Bach et al., 2015). Images hold the appeal of being easily interpretable by humans, making it simpler to validate model decisions by visual explanations. As such, many XAI methods provide explanations in the form of a relevance map over the input pixels. Time series do not benefit from the same properties, since they are by nature more difficult to understand (Rojat et al., 2021). While many time series models may be trained on raw time series data, the information (and thus relevance) may likely be found in a latent feature domain such as the frequency domain (Schröder et al., 2023). This challenge has only been addressed in (Vielhaben et al., 2024), where the authors use gradient-based methods to backpropagate the relevance into the frequency domain.
@@
Masking-based methods are an important tool in XAI due to their high performance across a multitude of domains (Petsiuk et al., 2019; Crabbé & Van Der Schaar, 2021; Highton et al., 2024). Masking-based approaches evolve around learning or estimating the relevace maps by iteratively applying binary masks to the input and measuring the resulting change in the output. Especially optimization-based masking approaches have shown great promise for time series data (Crabbé & Van Der Schaar, 2021; Enguehard, 2023; Liu et al., 2024). In optimization-based masking approaches, the relevance is learned by posing an objective that maximizes the number of masked out features in the input, while simultaneously minimizing the change in the model output (Fong et al., 2019). The mask is optimized via gradient descent through the model and the mask generating function. These methods can be difficult to use due to hyper-parameter tuning. Additionally, they suffer from the significant drawback of requiring access to the model gradients.
@@
Alternatively, model-agnostic methods require only access to the input and output of the network and are easily adaptable to a multitude of network architectures. Prominent model-agnostic methods are also based on masking and involve estimating the relevance via Monte Carlo sampling. This category includes RISE (Petsiuk et al., 2019) and RELAX (Wickstrøm et al., 2023). The sampling based methods have a few clear advantages: they do not require any complex optimization techniques, and we can directly constrain the sampling space to ensure compatibility with our data. While these methods have been successfully applied for time series (Mercier et al., 2022; Fedele et al., 2022), no previous work in this category has focused on providing relevances in another domain than the input domain.
@@
In this paper, we take inspiration from RISE and propose the first approach for masking-based XAI in the frequency domain of time series, FreqRISE. We:
@@
  • Propose FreqRISE, the first masking-based approach for providing relevance maps in the frequency domain.
@@
  • Provide a comprehensive evaluation of our proposed approach across several datasets and tasks.
@@ 2 Masking based explanations
Here, we present a new method for transforming masks between domains (Section 2.1) and use this new method to create FreqRISE (Section 2.2), the first masking-based XAI method operating in the frequency domain of time series.
@@ 2.1 Masking in a dual domain
While many time series models are trained directly on the raw time series data, the time domain is often insufficient for providing complete explanations (Schröder et al., 2023). This is due to most XAI methods being build on the assumption that the explanations are localized and sparse. However, if two classes are characterized by e.g. their frequency content, this information is neither localized nor sparse in the time domain. We, therefore, propose transforming the signals in to a domain where the information is assumed to be localized and sparse and to provide the explanations in this domain.
Our aim is to explain the black box model $f$ that outputs the class probabilities $\boldsymbol{y}\in\mathbb{R}^{C}$ from a time series input $\boldsymbol{X}\in\mathbb{R}^{V\times T}$, where $V$ is the number of input variables and $T$ is the length of the time series. Traditional masking-based methods use masks $\boldsymbol{M}\in[0,1]^{V\times T}$, to mask out features in the input space through elementwise multiplication, $\hat{\boldsymbol{X}}=\boldsymbol{X}\odot\boldsymbol{M}$ and observe the resulting changes in the model output:
$\hat{\boldsymbol{y}}=f(\hat{\boldsymbol{X}}).$ (1)
The result is therefore a relevance map in the input domain.
Instead, assume an invertible mapping into the domain of interest, $g:\mathbf{X}^{T}\rightarrow\mathbf{X}^{S}$. We can formulate an alternative masking-based approach, where the masks are applied in the new domain, $S$, after which the input is transformed back into the time domain, $T$. One example of such mapping is the discrete Fourier transform (DFT), which maps the signal into the frequency domain, $\mathbf{X}^{S}\in\mathbb{C}^{V\times F}$. We can then apply masks $\boldsymbol{M}^{S}\in[0,1]^{V\times F}$ in the frequency domain obtain the outputs through the inverse mapping:
$\hat{\boldsymbol{X}}=g^{-1}\left(g\left(\boldsymbol{X}\right)\odot\boldsymbol{M}^{S}\right).$ (2)
@@ 2.2 FreqRISE
A prominent masking-based framework is RISE (Petsiuk et al., 2019). RISE assumes that the masks $M$ are sampled from a distribution $D$. RISE then estimates the relevance, $R_{c,v,t}$, of class $c$ for a point in input space, $X_{v,t}$, as the expected value of $\hat{\boldsymbol{y}}_{c}$ (obtained using (1)) conditioned on $M_{v,t}=1$, i.e. that the point is observed:
$\boldsymbol{R}=\mathbb{E}_{M}[\hat{\boldsymbol{y}}\cdot\boldsymbol{M}].$ (3)
For a high accuracy classifier, we expect $\hat{\boldsymbol{y}}$ to be high when important time points are not masked out. In practice, we can estimate the relevance using Monte Carlo sampling. We produce $N$ masks and estimate the expected value as a weighted sum, normalized by the expectation over the masks:
$\hat{\boldsymbol{R}}=\frac{1}{N\cdot\mathbb{E}[M]}\sum_{i}^{N}\hat{\boldsymbol{y}}\cdot\boldsymbol{M}_{i}.$ (4)
In this work, we deal with univariate time series, i.e. $V=1$, however the methods can be extended to multivariate cases.
We combine the RISE framework with our proposed frequency and time-frequency masking and call our method FreqRISE.
@@ 3 Experimental setup
We conduct experiments on two datasets: the synthetic dataset presented in (Vielhaben et al., 2024) and AudioMNIST (Becker et al., 2024).
@@ 3.1 Synthetic data
We use the same synthetic dataset as in (Vielhaben et al., 2024). The dataset is specifically designed to have the salient information in the frequency domain and therefore allows to test the localization ability of the XAI methods. Each datapoint is created as a sum over $M$ sinusoids, where $M\sim\mathcal{U}(10,50)$ :
$X_{t}=\sum_{j}^{M}a_{j}\sin\left(\frac{2\pi t}{Tk_{j}}+\psi_{j}\right)+\sigma\epsilon.$ (5)
We set all $a_{j}=1$, randomly sample the phase $\psi_{j}\sim\mathcal{U}(0,2\pi)$, and add noise $\epsilon\sim\mathcal{N}(0,1)$ with standard deviation $\sigma$. The length of the signals is set to $T=2560$ and all frequency components are sampled as integer values within the range $k_{j}\sim\mathcal{U}\{1,59\}$. We train the models to detect a combination of frequencies, $k^{*}\in\{5,16,32,53\}$ from the time series signal. The classes are created from the powerset of $k^{*}$ and we therefore have 16 classes. Figure 1 shows a sample from the dataset.
@@
We train a 4-layer multilayer perceptron (MLP) with a hidden size of 64 on two different versions of the raw time series: one with noise level $\sigma=0.01$ and one with $\sigma=0.8$. In both cases, we use $10^{5}$ samples for training. We test the models on a test set of size $1000$ with no noise. Both models achieve an accuracy of 100%.
We use FreqRISE to compute relevance maps. We use the one-sided DFT to transform the signals to the frequency domain and sample binary masks, zeroing out single frequencies with $p=0.5$ from a Bernoulli distribution. We use $N=3000$ masks to obtain the relevance maps.
@@ 3.2 AudioMNIST
AudioMNIST is a dataset consisting of 30,000 audio recordings of spoken digits (0-9) repeated 50 times for each of the 60 speakers (12 female/48 male) (Becker et al., 2024). We follow (Becker et al., 2024) and downsample all signals to 8kHz and zero-pad all windows to 8000 samples.
We use the same 1D convolutional architecture as in (Becker et al., 2024) on the raw time series and train two versions: one trained for predicting the spoken digit and one trained to predict the gender of the speaker. We expect the salient information for the gender task and digit to be localized in the frequency domain and time-frequency domain, respectively. We obtain an accuracy of 96.9% on the digit classification and 98.6% for the gender task.
We use FreqRISE to compute relevance maps in the frequency and time-frequency domains and standard RISE in the time domain. We use 1000 data points from the gender and digit test sets respectively for testing the explanation methods. All relevance maps for AudioMNIST are estimated using $N=20,000$ samples. All masks are sampled from a Bernoulli distribution on a lower dimensional grid and linearly interpolated to create smooth masks. When computing the (Freq)RISE relevance maps, we use the logits prior to the softmax activation, since these show higher sensitivity to changes in the input.
We use the one-sided DFT to transform the signals to the frequency domain. We sample binary masks of size $200$ with $p=0.5$ and interpolate to size $F=4001$. In the time-frequency domain, we use the one-sided STDFT with a Hanning window of size 455 and an overlap of 420 samples between subsequent windows following (Becker et al., 2024). The binary masks are sampled as grids of size $25\times 25$ with $p=0.5$. Finally, in the time domain, we sample binary masks of size $200$ with $p=0.5$ and interpolate to size $T=8000$.
@@ 3.3 Baselines
As baselines, we use Integrated Gradients (IG) (Sundararajan et al., 2017) and Layer-wise Relevance Propagation (LRP) (Bach et al., 2015). When computing relevance maps in the frequency and time-frequency domains, we follow (Vielhaben et al., 2024) and equip the models with virtual inspection layers, which map the input into the relevant domains. Due to relevance conservation, we are restricted to using rectangular, non-overlapping windows for the STDFT. We therefore use rectangular windows of size 455 with no overlap.
@@ 3.4 Quantitative evaluation
Due to the lack of ground truth explanations (Hedström et al., 2023), quantitative evaluation for XAI is performed by measuring desirable properties (Hedström et al., 2023). Below we describe three such desirable properties that we use to quantify the quality of the relevance maps.
Localization: Localization scores are widely used to evaluate if explanations are co-located with a region of interest (Wickstrøm et al., 2023; Crabbé & Van Der Schaar, 2021). For the synthetic data, we know the ground truth explanations and can therefore use localization metrics to evaluate the methods. We use the relevance rank accuracy (Arras et al., 2022). Assuming a ground truth mask, $GT$, of size $K$ and a relevance map $\hat{\boldsymbol{R}}$, we take the $K$ points with the highest relevance. We then count how many of these values coincide with the positions in the ground truth mask. Formally, for $P_{topK}=\{p_{1},p_{2},\dots,p_{K}|\hat{R}_{p_{1}}>\hat{R}_{p_{2}}>\dots>\hat{R}_{p_{K}}\}$, i.e. $p$ denotes the position, we compute:
@@
Faithfulness: Faithfulness measures to what degree an explanation follows the predictive behaviour of the model and is a widely used measure for quantifying quality of explanations (Vielhaben et al., 2024; Liu et al., 2024). We follow prior works (Vielhaben et al., 2024; Crabbé & Van Der Schaar, 2021) and compute faithfulness as follows. Given a relevance map, $\hat{\boldsymbol{R}}$, we iteratively remove the $5\%,10\%,\dots,95\%$ most important features by setting the signal value to 0. We then evaluate the model performance as the mean probability of the true class and plot it to produce deletion plots. Finally, we compute the area-under-the-curve to produce our final faithfulness metric.
Complexity: Finally, we evaluate the complexity of the explanation as an estimate for the informativeness (Bhatt et al., 2021). The complexity is estimated as the Shannon entropy of the relevance maps.

### v2

# FreqRISE: Explaining time series using frequency masking

@@ Abstract
Time-series data are fundamentally important for many critical domains such as healthcare, finance, and climate, where explainable models are necessary for safe automated decision making. To develop explainable artificial intelligence in these domains therefore implies explaining salient information in the time series. Current methods for obtaining saliency maps assume localized information in the raw input space. In this paper, we argue that the salient information of a number of time series is more likely to be localized in the frequency domain. We propose FreqRISE, which uses masking-based methods to produce explanations in the frequency and time-frequency domain, and outperforms strong baselines across a number of tasks. The source code is available here: https://github.com/theabrusch/FreqRISE
@@ 1 Introduction
With the increasing development of systems for automated decision-making based on time series in critical domains such as healthcare [Phan_2022, 10285993], finance [GIUDICI2023104088] and climate forecasting [gonz2023a], the demand for safe and trustworthy machine learning models is ever rising. High accuracy and explainability are both necessary components in achieving safety and trustworthiness. Deep learning models have become a popular choice to obtain high accuracy for time series tasks [10.1145/3649448], but due to their complex reasoning process they are difficult to interpret. Explainable artificial intelligence (XAI) aims to open up the black box [samek2019explainable].
@@
Most advancements in the field of XAI have been made in explaining image data [tcav, rise, bach2015a]. Images hold the appeal of being easily interpretable by humans, making it simpler to validate model decisions by visual explanations. As such, many XAI methods provide explanations in the form of a relevance map over the pixels in the raw image [bach2015a, rise]. Time series do not benefit from the same properties, since they are by nature more difficult to understand [rojat2021a]. While many time series models may be trained on raw time series data, the information (and thus relevance) may likely be found in a latent feature domain such as the frequency domain [schroeder2023a]. This challenge has only been addressed in [vilrp], where the authors use gradient-based methods to backpropagate the relevance into the frequency domain.
@@
Masking-based methods are an important tool in XAI due to their high performance across a multitude of domains [rise, dynamask, 10.1007/978-981-97-1335-6_4]. Masking-based approaches learn or estimate the relevance maps by iteratively applying binary masks to the input and measuring the resulting change in the output. Especially optimization-based masking approaches have shown great promise for time series data [dynamask, extrmask, contralsp]. In optimization-based masking approaches, the relevance is learned by posing an objective that maximizes the number of masked out features in the input, while simultaneously minimizing the change in the model output [fong2019a]. The mask is optimized via gradient descent through the model and the mask generating function. However, these methods can be difficult to use due to hyper-parameter tuning. Additionally, they suffer from the significant drawback of requiring access to the model gradients.
@@
Alternatively, model-agnostic methods require only access to the input and output of the network and are easily adaptable to a multitude of network architectures. Prominent model-agnostic methods are also based on masking and involve estimating the relevance via Monte Carlo sampling. This category includes RISE [rise] and RELAX [relax]. The sampling based methods have a few clear advantages: they do not require any complex optimization techniques, and we can directly constrain the sampling space to ensure compatibility with our data. While these methods have been successfully applied for time series [mercier2022a, riseaudio], no previous work in this category has focused on providing relevance maps in another domain than the input domain.
@@
In this paper, we take inspiration from RISE and propose the first approach for masking-based XAI in the frequency domain of time series, FreqRISE.
We:
@@
  • Propose FreqRISE, the first masking-based approach for providing relevance maps in the frequency and time-frequency domain.
@@
  • Provide a comprehensive evaluation of our proposed approach across two datasets and four tasks.
@@ 2 Masking based explanations
Figure 1: The FreqRISE framework shown when, as an example, using the Short-Time Discrete Fourier Transform (STDFT) as the invertible mapping, $g$. $g$ maps $\mathbf{X}^{T}$ into the domain of interest, $g:\mathbf{X}^{T}\rightarrow\mathbf{X}^{S}$. Here, we sample $N$ masks and produce $N$ masked versions of $\mathbf{X}^{S}$. Using the inverse of $g$, we map back to the time domain, obtain $\hat{y}_{c}$, and compute the relevance map as a weighted average over the masks.
Here, we present a new method for transforming masks between domains (Section 2.1) and use this new method to create FreqRISE (Section 2.2), the, to our knowledge, first masking-based XAI method operating in the frequency domain of time series.
@@ 2.1 Masking in a dual domain
While many time series models are trained directly on the raw time series data, the time domain is often insufficient for providing complete explanations [schroeder2023a]. This is due to most XAI methods being built on the assumption that the explanations are localized and sparse. However, if two classes are characterized by e.g. their frequency content, this information is neither localized nor sparse in the time domain. We, therefore, propose transforming the signals to a domain where the information is assumed to be localized and sparse and to provide the explanations in this domain.
Our aim is to explain the black box model $f(\cdot)$ that outputs the class probabilities $\boldsymbol{y}\in\mathbb{R}^{C}$ from a time series input $\boldsymbol{X}\in\mathbb{R}^{V\times T}$, where $V$ is the number of input variables and $T$ is the length of the time series. Masking-based methods, such as RISE [rise] and RELAX [relax], use masks $\boldsymbol{M}\in\{0,1\}^{V\times T}$ sampled from distribution $\mathcal{D}$, to mask out features in the input space. Here the $n$ index identifies the $n$-th sampled mask. The masking is done through elementwise multiplication such that $\hat{\boldsymbol{X}}\left(\boldsymbol{M}\right)=\boldsymbol{X}\odot\boldsymbol{M}$. The resulting change in the model output can then be computed:
$\hat{\boldsymbol{y}}\left(\boldsymbol{M}\right)=f(\hat{\boldsymbol{X}}\left(\boldsymbol{M}\right)).$ (1)
The result is therefore a relevance map over the input features in the input domain.
Instead, assume an invertible mapping from the time domain, $T$, into the domain of interest, $S$, $g:\mathbf{X}^{T}\rightarrow\mathbf{X}^{S}$. We can formulate an alternative masking-based approach, where the masks are applied in the new domain after which the input is transformed back into the time domain. One example of such mapping is the discrete Fourier transform (DFT), which maps the signal into the frequency domain, $\mathbf{X}^{S}\in\mathbb{C}^{V\times F}$. We can then apply masks $\boldsymbol{M}\in\{0,1\}^{V\times F}$ in the frequency domain and obtain the masked input in the time domain via the inverse mapping:
$\hat{\boldsymbol{X}}\left(\boldsymbol{M}\right)=g^{-1}\left(g\left(\boldsymbol{X}\right)\odot\boldsymbol{M}\right).$ (2)
@@ 2.2 FreqRISE
A prominent masking-based framework is RISE [rise]. RISE assumes that the masks $\boldsymbol{M}$ are sampled from a distribution $\mathcal{D}$. RISE then estimates the relevance, $\boldsymbol{R}_{c}\left(\lambda\right)$, of class $c$ for a point in input space, $\lambda$, as the expected value of the class probability $\hat{y}_{c}(\boldsymbol{M})$ (obtained using (1)) under the distribution of $\boldsymbol{M}$ conditioned on $\boldsymbol{M}\left(\lambda\right)=1$, i.e. that the point is observed. Here, according to our definition, $\lambda=(v,t)$ if defined in the time domain, $\lambda=(v,f)$ in the frequency domain, and $\lambda=(v,t,f)$ in the time-frequency domain. For a full derivation of the mathematical details, see the appendix.
For a high accuracy classifier, we expect $\hat{y}_{c}$ to be to be low when the most important features are removed and, reversely, to be high when important features are not masked out. In practice, we can estimate the relevance using Monte Carlo sampling. We produce $N$ masks and estimate the expected value as a weighted sum, normalized by the expectation over the masks [rise]:
$\hat{\boldsymbol{R}}_{c}=\frac{1}{N\cdot\mathbb{E}[\boldsymbol{M}]}\sum_{n=1}^{N}\hat{y}_{c}(\boldsymbol{M}_{n})\cdot\boldsymbol{M}_{n}.$ (3)
Since the masks are applied in the dual domain, the resulting relevance map is a map over relevances in this dual domain. In this work, we deal with univariate time series, i.e. $V=1$, however the methods can be extended to multivariate cases.
We combine the RISE framework with our proposed frequency and time-frequency masking and call our method FreqRISE. The FreqRISE framework is shown in Figure 1 when using STDFT as the invertible mapping.
@@ 3 Experimental setup
We conduct experiments on two datasets: the synthetic dataset presented in [vilrp] and AudioMNIST [audiomnist]. Both datasets have been shown to have salient information in either the frequency or time-frequency domain [vilrp].
@@ 3.1 Synthetic data
We use the same synthetic dataset as in [vilrp]. The dataset is specifically designed to have the salient information in the frequency domain and therefore allows us to test the localization ability of the XAI methods. Each data point is created as a sum over $J$ sinusoids. $J$ is sampled uniformly from $\mathcal{U}(10,50)$ :
$X_{t}=\sum_{j}^{J}a_{j}\sin\left(\frac{2\pi t}{Tk_{j}}+\psi_{j}\right)+\epsilon.$ (4)
We set all $a_{j}=1$, randomly sample the phase $\psi_{j}\sim\mathcal{U}(0,2\pi)$, and add normally distributed noise $\epsilon\sim\mathcal{N}(0,\sigma)$. The length of the signals is set to $T=2560$ and all frequency components are sampled as integer values within the range $k_{j}\sim\mathcal{U}\{1,59\}$. We train the models to detect a combination of frequencies, $k^{*}\in\{5,16,32,53\}$ from the time series signal. The classes are created from the powerset of $k^{*}$, and therefore we have 16 classes. Figure 2 shows a sample from the dataset.
@@
We train a 4-layer multilayer perceptron (MLP) with a hidden size of 64 on two different versions of the raw time series: one with noise level $\sigma=0.01$ and one with $\sigma=0.8$. In both cases, we use $10^{5}$ samples for training. We tested the models on a test set of size $1000$ without noise. The two models trained on the synthetic dataset both achieve an accuracy of 100%.
We use FreqRISE to compute relevance maps. We use the one-sided DFT to transform the signals to the frequency domain and sample binary masks, zeroing out single frequencies with $p=0.5$ from a Bernoulli distribution. Similar to recent work [relax], we use $N=3000$ masks to obtain relevance maps.
@@ 3.2 AudioMNIST
AudioMNIST is a dataset consisting of 30,000 audio recordings of spoken digits (0-9) repeated 50 times for each of the 60 speakers (12 female/48 male) [audiomnist]. We follow [audiomnist] and downsample all signals to 8kHz and zero-pad all windows to 8000 samples.
We use the same 1D convolutional architecture as in [audiomnist] on the raw time series and train two versions: one for predicting the spoken digit and one to predict the gender of the speaker. The model trained on the digit classification task achieves and accuracy of 96.9%, while the model trained on the gender tasks achieves an accuracy of 98.6%.
Since the fundamental frequency is known to be discriminative for gender [baken2000clinical], we expect the salient information for the gender task to be localized in the frequency domain. The digit task, however, will likely be localized in both time and frequency, since we expect the numbers to be discriminated both through their formant information [brend1990a] (frequency domain) and the ordering of these in time (time domain).
We use FreqRISE to compute relevance maps in the frequency and time-frequency domains and standard RISE in the time domain. We use 1000 data points from the gender and digit test sets respectively for testing the explanation methods. All relevance maps for AudioMNIST are estimated using $N=10,000$ sampled masks. All masks are sampled from a Bernoulli distribution with $p=0.5$ on a lower dimensional grid and linearly interpolated to create smooth masks. When computing the (Freq)RISE relevance maps, we use the logits prior to the softmax activation, since these show higher sensitivity to changes in the input.
In the time-frequency domain, we use the one-sided STDFT with a Hanning window of size 455 and an overlap of 420 samples between subsequent windows following [audiomnist]. In STDFT domain, the binary masks are sampled as grids of size $25\times 25$. In the time and frequency domains, we use grids of size $200$.
@@ 3.3 Baselines
As baselines, we use Integrated Gradients (IG) [pmlr-v70-sundararajan17a] and Layer-wise Relevance Propagation (LRP) [bach2015a]. When computing relevance maps in the frequency and time-frequency domains, we follow [vilrp] and equip the models with virtual inspection layers, which map the input into the relevant domains. Due to relevance conservation, we are restricted to using rectangular, non-overlapping windows for the STDFT. We therefore use rectangular windows of size 455 with no overlap. IG and LRP assign both negative and positive values to the relevance maps. Here, we consider only positive relevance when evaluating relevance maps.
@@ 3.4 Quantitative evaluation
Due to the lack of ground truth explanations [hedstrom2023metaquantus], quantitative evaluation for XAI is performed by measuring desirable properties [hedstrom2023quantus]. In the following, we describe three such desirable properties that we use to quantify the quality of the relevance maps.
Localization: Localization scores are widely used to evaluate if explanations are colocated with a region of interest [relax, dynamask]. For the synthetic data, we know the ground truth explanations and can therefore use localization metrics to evaluate the methods. We use the relevance rank accuracy [clevr-xai]. Assuming a ground truth mask, $GT$, of size $K$ and a relevance map $\hat{\boldsymbol{R}}_{c}$, we take the $K$ points with the highest relevance. We then count how many of these values coincide with the positions in the ground truth mask. Formally, for $P_{topK}=\{p_{1},p_{2},\dots,p_{K}|\hat{R}_{p_{1}}>\hat{R}_{p_{2}}>\dots>\hat{R}_{p_{K}}\}$, i.e. $p$ denotes the position, we compute:
@@
Faithfulness: Faithfulness measures to what degree an explanation follows the predictive behaviour of the model and is a widely used measure for quantifying quality of explanations [vilrp, contralsp]. We follow prior works [vilrp, dynamask] and compute faithfulness as follows. Given a relevance map, $\hat{\boldsymbol{R}}_{c}$, we iteratively remove the $5\%,10\%,\dots,50\%$ most important features by setting the signal value to 0. We then evaluate the model performance as the mean probability of the true class and plot it to produce deletion plots. Finally, we compute the area-under-the-curve (AUC) to produce our final faithfulness metric. A low AUC means that the explanation is faithful to the model.
Complexity: Finally, we evaluate the complexity of the explanation as an estimate for the informativeness [complexity]. The complexity is estimated as the Shannon entropy of the relevance maps.
3.5 Post-processing of relevance maps
Generally, FreqRISE (as well as other sampling-based explainability methods) suffer from the limitation that they cannot assign 0 relevance to any part of the input. This is due to the weighted average computation of the relevance map which means that the baseline relevance given to unimportant areas will fluctuate around some lower value (typically close to zero). This behaviour increases the complexity of the relevance maps. Therefore, we explore post-processing of the relevance maps as a mean to decrease the complexity. Given a relevance map $\hat{\boldsymbol{R}}_{c}$, we produce a post-processed relevance map $\hat{\boldsymbol{R}}_{c}^{p}$ by computing the $p$-th quantile, $q_{p}$ and setting all elements, $\lambda$, with value below $q_{p}$ to 0:
$\hat{\boldsymbol{R}}_{c}^{p}\left(\lambda\right)=\begin{cases}\hat{\boldsymbol{R}}_{c}^{p}\left(\lambda\right),&\text{if $\hat{\boldsymbol{R}}_{c}^{p}\left(\lambda\right)\geq q_{p}$}.\\ 0,&\text{otherwise}.\end{cases}$ (6)
We compare the faithfulness and complexity trade-off for post-processed relevance maps of FreqRISE with different thresholds $p$ on the AudioMNIST dataset.